Q1. An argon-ion laser emits a blue-green beam of light with a wavelength of 488nm in a vacuum. Compute the energy in joules for the two energy states for the atomic transition which produces this light?
1
Expert's answer
2020-06-08T10:16:44-0400

The energy of the photon will be


"E=\frac{hc}{\lambda}=\frac{(6.626\cdot10^{-34})(3\cdot10^8)}{488\cdot10^{-9}}=4.047\cdot10^{-19}\text{ J},"

or 2.54 eV.